An exclusive pen manufacturer sells 4,000 pens per month at a price of £40 each. When the price is reduced to £30 sales increase to 6,000 pens per month. (a) Calculate the price elasticity of demand for the pens over this price range. (b) Is demand elastic, unit elastic or inelastic? (c) Calculate the change in revenue due to the change in price


1
Expert's answer
2021-01-21T07:58:31-0500

"Pe=(Qd2-qd1\/q1)\/(p2-p1\/p1)"

=(6000-4000)/ 4000=0.5

=30-40/40=-0.25

=0.5/-0.25=-2

=3448

(B) the demand is inelastic

(C) total revenue="Qd2*P2)-(Q1*P1)"

(6000*30)-(4000*40)

=2000